OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12

OpenStudy (zzr0ck3r):

isolate x first -x-2y = -4 -x = -4+2y x=4-2y